OpenStudy (anonymous):

ASSIGNMENT HELP PLEASE! 7 QUESTIONS WILL MEDAL AND FAN!

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (03.07) Is the line for the equation x = 4 horizontal or vertical? What is the slope of the line? Select the best answer. The line is horizontal. The slope is 0. The line is vertical. The slope is undefined. The line is horizontal. The slope is undefined. The line is vertical. The slope is 0. Question 2 (Multiple Choice Worth 2 points) (03.07) Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14 Question 3 (Multiple Choice Worth 2 points) (03.07) Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7 Question 4 (Multiple Choice Worth 1 points) (03.07) Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1 Question 5 (Multiple Choice Worth 2 points) (03.07) Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = -one halfx + 3 y = -one halfx + 6 y = 2x + 3 Question 6 (Multiple Choice Worth 1 points) (03.07) Graph with a line going through points zero comma two point five and four comma two point five. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 Question 7 (Multiple Choice Worth 1 points) (03.07) Select the coordinates of two points that are on the line x = 5. (5, 0) and (0, 5) (5, 0) and (5, -2) (5, 0) and (0, 2) (0, 5) and (0, 2)

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

please help me! @iGreen

OpenStudy (anonymous):

1. the line is vertical with a slope of 0.

OpenStudy (anonymous):

im terrible at stuff like this thats all i know sorry :(